OpenStudy (anonymous):

Solve by Factoring: x2 + 5x - 6 = 0 What are the solutions for x? x = 2, x = 3 x = 6, x = -1 x = -6, x = 1 none of the above

OpenStudy (anonymous):

Its the third option.

OpenStudy (anonymous):

-6,1

hero (hero):

x^2 +5x - 6 = 0 x^2 + 6x-1x -6 = 0 x(x+6)-1(x+6) = 0 (x+6)(x-1) = 0 x = -6,1